OpenStudy (anonymous):

tell whether the ordered pair ( 5,-3) is a solution of the system { -3x+2y=-21} {-x-y=-2}

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@Peaches15

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@LynFran @demiac259 @Destinyyyy @DullJackel09 @lostinrome @ChrystalC @Chris01

OpenStudy (texaschic101):

you can sub in your points into each equation and if the equations come out true (both of them), then it is a solution. Understand ?

OpenStudy (texaschic101):

Let me set up the first equation... -3x + 2y = -21 (x = 5 and y = -3) -3(5) + 2(-3) = -21 now work that out and tell me if it is true.

OpenStudy (texaschic101):

do the same for the other equation... -x - y = -2 -5 - (-3) = -2 Is that true ? If they are BOTH true, then those points are a solution, But if even 1 of them is false, then it is not a solution

OpenStudy (texaschic101):

Let me know what you come up with :)

OpenStudy (anonymous):

-5-(-3)=-2 @texaschic101

OpenStudy (texaschic101):

-3x + 2y = -21 -3(5) + 2(-3) = -21 -15 - 6 = -21 -21 = -21 (true) -x - y = -2 -5 - (-3) = -2 -5 + 3 = -2 -2 = -2 (true) so if they are both true, which they are, then the solution is true and (5,-3) are points that satisfy this set of equations :)